(A) Aluminum chloride reacts with sodium hydroxide to produce an aluminum hydroxide precipitate. Write a balance chemical reaction.
(B) How many moles aluminum hydroxide and sodium chloride would be produced if 2.0 molesof sodium hydroxide are reacted with excess aluminum chloride?
1
Expert's answer
2020-06-08T15:41:26-0400

Solution:

Part A:

Aluminum chloride react with sodium hydroxide to produce aluminium hydroxide and sodium chloride.

The balanced chemical equation:

AlCl3(aq) + 3NaOH(aq) → Al(OH)3(s) + 3NaCl(aq)


Ionic equation: Al3+(aq) + 3Cl-(aq) + 3Na+(aq) + 3OH-(aq) → Al(OH)3(s) + 3Na+(aq) + 3Cl-(aq)

Net ionic equation: Al3+(aq) + 3OH-(aq) → Al(OH)3(s)


Part B:

The balanced chemical equation:

AlCl3(aq) + 3NaOH(aq) → Al(OH)3(s) + 3NaCl(aq)

According to the chemical equation:

1) n(NaOH) = n(NaCl) = 2.0 mol

Moles of NaCl = 2.0 mol

2) n(NaOH)/3 = n(Al(OH)3)

Moles of Al(OH)3 = Moles of NaOH / 3 = (2.0 mol) / 3 = 0.67 mol


0.67 mol of aluminum hydroxide (Al(OH)3) and 2.0 mol of sodium chloride (NaCl) would be produced.